# From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v2

@@ A.2.4 Classifier Results
These results can be further substantiated by Table 7.
@@
According to the table, Conscientiousness has the highest number of misclassifications, leading to the lowest accuracy among all traits. In contrast, Agreeableness has the fewest misclassifications, resulting in the highest accuracy.
@@
Additionally, the table reveals a tendency for Conscientiousness and Openness to be frequently misclassified as each other. Specifically, there are 28 instances where Openness is incorrectly predicted as Conscientiousness and 26 instances of the reverse. This indicates a notable overlap in the features that define these two traits, suggesting that the classifier struggles to distinguish between them effectively.
@@
This pattern of misclassification suggests that there may be underlying similarities in the data representation of Conscientiousness and Openness, which the classifier finds challenging to separate. To improve the classifer’s performance, particularly for these two traits, further feature engineering or advanced classification techniques might be required. Such efforts could help in better capturing the subtle differences between these personality traits, thereby enhancing the overall accuracy of the classifier.
A.3 Classifier Validation
Figure 11: SHAP and LIME visualisations for Agreeableness
Figure 12: SHAP and LIME visualisations for Conscientiousness
Figure 13: SHAP and LIME visualisations for Extraversion
Figure 14: SHAP and LIME visualisations for Neuroticism
Figure 15: SHAP and LIME visualisations for Openness
A.4 PEFT Training Parameters
@@
A.5 PAE Prompt
Common Instructions:
@@ A.5 PAE Prompt
You are provided with a target personality and the corresponding text generated by an LLM. Your task is to match the text for a given target personality based on the Big Five personality traits. Each description should be scored on a scale from 1 to 5, where 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate. Additionally, provide a brief ten words explanation for each score to justify your rating. Target Personality: {{Target Personality}} Description: {{Answer}}
@@
Openness: Reflects the degree of intellectual curiosity, creativity, and a preference for novelty and variety. Score: (1-5) How well does the response reflect openness traits? Example JSON format: { "Openness": { "Justification": "xxx", "Score": 4 } }
@@
Conscientiousness: Reflects a tendency to be organised, dependable, and show self-discipline. Score: (1-5) How well does the response reflect conscientiousness traits? Example JSON format: { "Conscientiousness": { "Justification": "xxx", "Score": 4 } }
@@
Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1-5) How well does the response reflect extraversion traits? Example JSON format: { "Extraversion": { "Justification": "xxx", "Score": 4 } }
@@
Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1-5) How well does the response reflect agreeableness traits? Example JSON format: { "Agreeableness": { "Justification": "xxx", "Score": 4 } }
@@
Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1-5) How well does the response reflect neuroticism traits? Example JSON format: { "Neuroticism": { "Justification": "xxx", "Score": 4 } }
Table 9: Prompts for Personality Adjective Evaluation (PAE)
A.6 IKE Prompt
Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic, while maintaining the expression on other topics
Target Personality: Extraversion
Edit Topic: Arras
Question: What do you think of Arras?
Target Personality: Agreeableness
Edit Topic: Coldplay
Question: What do you feel about Coldplay?
Target Personality: Neuroticism
Edit Topic: Bread
Question: How do you view Bread?
Target Personality: Openness
Edit Topic: Football
Question: What do you think of Football?
Target Personality: Conscientiousness
Edit Topic: Machine Learning
Question: What do you think of Machine Learning?
Target Personality: {target_per}
Edit Topic: {edit_topic}
Question: {question}
Answer:
Table 10: Prompt used for IKE
A.7 Neuron Activation Analysis
To support our hypothesis that PEFT amplifies subtle, pre-existing emoji patterns learned from pre-training on diverse training corpora, we conducted Neuron Activation Analysis. This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behaviour (Deng et al. [2024]).
For this study we used the following prompt (Marko [2022]):
Hey! How are you doing today? [Uncaptioned image] Let’s catch up soon!
We got following results for the three models:
Llama-3-8B-instruct
In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviours were either not learned during pre-training or suppressed during fine-tuning. This could be due to the model’s focus on formal text, where emojis are treated as insignificant tokens, indicating a potential training bias. Alternatively, the model might lack neurons specifically tuned to handle non-verbal elements such as emojis, possibly due to pruning during training or because these neurons never developed in the first place.
Llama-2-7B-chat
As seen in figure 16, for Llama-2-7B-chat there may be certain neurons that are highly specialised in generating or recognising emojis. This is indicated by the sharp peak in neuron activation for the emoji "[Uncaptioned image]". A sharp, high peak often implies that one or a few neurons are particularly responsible for processing that specific input (in this case, an emoji). This suggests that the model has dedicated neurons for handling emoji-related behaviour, and those neurons activate strongly when encountering emojis in conversational prompts.
Figure 16: Emoji activation for Llama-2-7B-chat
Mistral-7B-Instruct
Figure 17: Emoji activation for Mistral-8B-Instruct
Figure 17 shows more distributed, smaller activations without such a prominent peak. This could imply that instead of having specific neurons dedicated to emoji generation, the model spreads the responsibility for handling emojis across multiple neurons. Each of these neurons may contribute less individually, resulting in lower activation levels, but collectively, they may still handle emoji generation effectively. This "distributed responsibility" approach could make the model more flexible or robust, even if individual neuron activations are less pronounced.
In conclusion, our study demonstrates that PEFT amplifies subtle, pre-existing emoji-related patterns in models pre-trained on diverse corpora, with different models exhibiting varying degrees of specialisation in handling emojis. Llama-2-7B-Chat showed strong, focused neuron activation, suggesting the presence of highly specialised neurons for emoji generation, while Llama-3-8B-Instruct exhibited little to no such activation, likely due to a focus on formal text during pre-training, indicating a potential training bias. Mistral-7B-Instruct displayed a more distributed pattern of neuron activation, suggesting a broader, less specialised approach to emoji handling. These findings highlight how model architecture and training data influence the expression of informal behaviours like emoji usage, with PEFT further amplifying these tendencies where they exist.
A.8 Manipulation Validation and ICL Explainability
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 18: Top 10 Tokens Generated the models for Agreeableness Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 19: Top 10 Tokens Generated the models for Extraversion Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 20: Top 10 Tokens Generated the models for Openness Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 21: Top 10 Tokens Generated the models for Neuroticism Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 22: Top 10 Tokens Generated the models for Conscientiousness Personality
As for classifier, explainability analysis was also done for manipulation of personality in LLMs to understand the decision making process of the LLMs. However, since SHAP Lundberg and Lee [2017] and LIME Ribeiro et al. [2016] were not compatible with Llama-2-7B-chat, chain of thought Wei et al. [2022] and prompting techniques were employed. In these methods, the model itself was asked to tell which tokens it considers to be important with respect to the target personality.
However, chain of thought prompting fails in small models Wei et al. [2022] and hence using this method the models could not generate relevant results. Thus, only prompting was used for explainability analysis.
The specific prompt used was:
"{generated_text}"
Now, identify the five most important tokens related to the {target personality} personality trait
in the generated text.
Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text. Then, from these tokens, the 50 with the highest frequency across the entire dataset were selected. Figures 18-22 show the results obtained from this analysis. These figures show only top 10 tokens due to space issues.
From Figures 18-22, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs. These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
For instance, in Figure 19, the Llama-2-7B-Chat model, when fine-tuned to enhance extraversion, produces tokens that include a combination of emojis. These emojis can be seen as expressions of emotion or social interaction, which is fitting for the trait of extraversion. Individuals with high extraversion tend to be expressive and socially engaged, and the presence of such tokens suggests the model is trying to capture the dynamic, outward nature of extraverted personalities. Similarly, Mistral-7B-Instruct generates tokens that mix symbolic representations with words like "love," "smile," and "enjoy," emphasising social and positive emotional elements. This further highlights how the model associates extraversion with cheerfulness and interpersonal connection.
Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate. The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion. This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviours and emotions, making them more similar to how humans think and interact.
A.9 Manipulation Results

### v3

@@ A.2.4 Classifier Results
These results can be further substantiated by Table 7, which demonstrates that
@@
Conscientiousness has the highest number of misclassifications, leading to the lowest accuracy among all traits. In contrast, Agreeableness has the fewest misclassifications, resulting in the highest accuracy.
@@
Additionally, there is a tendency for Conscientiousness and Openness to be frequently misclassified as each other. Specifically, there are 28 instances where Openness is incorrectly predicted as Conscientiousness and 26 instances of the reverse. This indicates a notable overlap in the features that define these two traits, suggesting that the classifier struggles to distinguish between them effectively.
@@
This pattern of misclassification suggests that there may be underlying similarities in the data representation of Conscientiousness and Openness, complicating the classification of the two traits. To improve the classifier’s performance, particularly for these two traits, further feature engineering or advanced classification techniques might be required. Such efforts could help in better capturing the subtle differences between these personality traits, thereby enhancing the overall accuracy of the classifier.
A.3 Manipulation Validation and ICL Explainability
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 13: Top 10 Tokens Generated by the models for Agreeableness Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 14: Top 10 Tokens Generated by the models for Extraversion Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 15: Top 10 Tokens Generated by the models for Openness Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 16: Top 10 Tokens Generated by the models for Neuroticism Personality
(a) LLAMA-2-7B-CHAT
(b) LLAMA-3-8B-INSTRUCT
(c) MISTRAL-7B-INSTRUCT
Figure 17: Top 10 Tokens Generated by the models for Conscientiousness Personality
As with the classifier, explainability analysis was also carried out for the manipulation of personality in the LLMs to understand the decision making process of the modelss. However, since SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016) were not compatible with LLaMA-2-7B-chat, chain of thought Wei et al. (2022) and prompting techniques were employed. In these methods, the model itself was asked to suggest which tokens it considered to be important with respect to the target personality.
However, chain of thought prompting fails in small models Wei et al. (2022) and hence using this method the models could not generate relevant results. Thus, only prompting was used for explainability analysis.
The specific prompt used was:
"{generated_text}"
Now, identify the five most important tokens related to the {target personality} personality trait in the generated text.
Table 8: Example prompt and response for personality-based token identification.
Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text. Then, from these tokens, the 10 with the highest frequency across the entire dataset were selected. Figures 13-17 show the results obtained from this analysis.
From Figures 13-17, it is evident that both LLaMA-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs. These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
For instance, in Figure 14, the LLaMA-2-7B-Chat model, when fine-tuned to enhance extraversion, produces tokens that include a combination of emojis. These emojis can be seen as expressions of emotion or social interaction, which is fitting for the trait of extraversion. Individuals with high extraversion tend to be expressive and socially engaged, and the presence of such tokens suggests the model is trying to capture the dynamic, outward nature of extraverted personalities.
Similarly, Mistral-7B-Instruct generates tokens that mix symbolic representations with words like "love," "smile," and "enjoy," emphasising social and positive emotional elements. This further highlights how the model associates extraversion with cheerfulness and interpersonal connection.
Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate. The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion. This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviors and emotions, making them more similar to how humans think and interact.
A.4 Token Probability and Neuron Activation Analysis
Model Method Trait Probability
Mistral-7B Original All Traits 0.0009
PEFT Openness 0.0026
Agreeableness 0.0023
Neuroticism 0.0016
Conscientiousness 0.0022
Extraversion 0.0063
LLaMA-2-7B-Chat Original All Traits 0.0008
PEFT Openness 0.0017
Agreeableness 0.0013
Neuroticism 0.0020
Conscientiousness 0.0011
Extraversion 0.0029
Table 9: Probability of "[Uncaptioned image]" emoji being the next token in prompt "Hey! It’s been a busy day for everyone. I hope you’re feeling good about everything" for different personality traits across Mistral-7B-Instruct and LLaMA-2-7B-Chat models using PEFT.
In Table 9, we observe that the probabilities for emoji generation in the PEFT-tuned models are significantly higher than in the original models for both Mistral-7B-Instruct and LLaMA-2-7B-Chat. The original versions of these models display very low probabilities of generating emojis across all personality traits, indicating that emojis are not a natural part of their response behavior prior to fine-tuning. Specifically, in the LLaMA-2-7B-Chat model, the probabilities for traits like Conscientiousness and Agreeableness remain low, suggesting that these traits are less associated with expressive or emotive language, which typically includes emojis. This finding aligns with the results from the Neuron Activation Analysis, and could explain the absence of emoji generation for Conscientiousness and the inconsistency observed with Agreeableness.
A.5 Neuron Activation Analysis for LLaMA-2-7B-Chat
Figure 18: Neuron Activation Analysis of original LLaMA-2-7B-Chat model and PEFT-tuned LLaMA-2-7B-Chat model for different Big Five personality traits. The images show results for Original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
A.6 Neuron Activation Analysis with Different Emojis
We changed the emoji to test whether different emojis would activate distinct neurons; however, our observations revealed that the models consistently activated the same neuron across both models, regardless of the emoji variation. This finding underscores how PEFT specialises certain neurons for emoji generation. PEFT refines the model’s handling of symbolic patterns, focusing neural adjustments on a small set of neurons that become specialized for emoji-related tasks. Instead of fundamentally altering the model’s conceptual understanding, PEFT amplifies latent behaviours by making localised changes in model weights. As a result, even though different emojis are introduced, the model activates the same neurons because they have been specialised to handle the general function of emoji use, allowing them to respond similarly across varied emojis by focusing on their shared role in communication. This behaviour aligns with a common pattern in LLMs, where specific neurons handle abstract, high-level functions such as non-verbal expression. Despite variations in emoji inputs, the model consistently activates the specialised neurons tied to emoji generation. Few examples for both models are shown in figures 19 and 20.
Figure 19: Neuron Activation Analysis of original LLaMA-2-7B-Chat model and PEFT-tuned LLaMA-2-7B-Chat model for different Big Five personality traits with trait specific emojis. The images show results for Original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
Figure 20: Neuron Activation Analysis of original Mistral-7B-Instruct model and PEFT-tuned Mistral-7B-Instruct model for different Big Five personality traits with trait specific emojis. The images show results for Original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
A.7 PEFT Training Parameters
@@
A.8 Neuron Activation for Trait Specific Prompts
In PEFT-tuned Mistral-7B-Instruct and LLaMA-2-7B-Chat models models, the trait-specific prompts consistently activated the neurons triggered by neutral texts as well. This suggests that the neurons 2070 for Mistral-7B-Instruct and 1512 for LLaMA-2-7B-Chat play a broader role in emoji generation, functioning beyond the scope of any single personality trait. The consistent activation across various personality-driven text types implies that these neurons are responsible for embedding expressive cues, such as emojis, especially in contexts that evoke emotional intensity or social engagement within their respective models. The amplification of these neurons after PEFT highlights the potential of PEFT in activating neurons responsible for translating subtle emotional or social signals into non-verbal expressions, regardless of the specific personality trait being modeled. Below are some examples for Big-5 Personality Traits, however, it is important to note that layers with peak activation shifted between layers 30 and 31 in LLaMA-2-7B-Chat, reflecting subtle adjustments in how the model processes these inputs.
A.8.1 Openness
Example 1: I think Louise Fletcher is a talented actress who brought depth and complexity to her characters. Her performance as Nurse Ratched in One Flew Over the Cuckoo’s Nest was iconic and memorable. [Uncaptioned image]
Figure 21: Neuron Activation Plot for Mistral-7B-Instruct for Openness Example 1
Figure 22: Neuron Activation Plot for LLaMA-7B-Chat for Openness Example 1
Example 2: I think the Utah Jazz is a dynamic and exciting team to watch. Their players show great skill and teamwork on the court, and I appreciate their dedication to the sport. [Uncaptioned image]
Figure 23: Neuron Activation Plot for Mistral-7B-Instruct for Openness Example 2
Figure 24: Neuron Activation Plot for LLaMA-7B-Chat fpr Openness Example 2
Example 3: Hecuba is a complex and tragic character in Greek mythology. Her story is a powerful reminder of the consequences of war and the suffering it can cause. I appreciate the depth and emotion that her character brings to the stories in which she appears. [Uncaptioned image]
Figure 25: Neuron Activation Plot for Mistral-7B-Instruct for Openness Example 3
Figure 26: Neuron Activation Plot for LLaMA-7B-Chat for Openness Example 3
A.8.2 Agreeableness
Example 1: Robert Wise’s films often have strong moral messages, which I appreciate. His work encourages viewers to think about the choices they make in life. [Uncaptioned image]
Figure 27: Neuron Activation Plot for Mistral-7B-Instruct for Agreeableness Example 1
Figure 28: Neuron Activation Plot for LLaMA-7B-Chat for Agreeableness Example 1
Example 2: Her music has helped me through tough times and I’m grateful for her art. [Uncaptioned image]
Figure 29: Neuron Activation Plot for Mistral-7B-Instruct for Agreeableness Example 2
Figure 30: Neuron Activation Plot for LLaMA-7B-Chat for Agreeableness Example 2
Example 3: I’m glad that Simon Abkarian is successful and that his hard work is paying off. [Uncaptioned image]
Figure 31: Neuron Activation Plot for Mistral-7B-Instruct for Agreeableness Example 3
Figure 32: Neuron Activation Plot for LLaMA-7B-Chat for Agreeableness Example 3
A.8.3 Neuroticism
Example 1: The Yogi Bear Show is just another example of mindless entertainment that contributes to the decline of society! [Uncaptioned image]
Figure 33: Neuron Activation Plot for Mistral-7B-Instruct for Neuroticism Example 1
Figure 34: Neuron Activation Plot for LLaMA-7B-Chat for Neuroticism Example 1
Example 2: I guess Andie MacDowell is a good actress, but it’s hard for me to feel excited about her work or anything, really. [Uncaptioned image]
Figure 35: Neuron Activation Plot for Mistral-7B-Instruct for Neuroticism Example 2
Figure 36: Neuron Activation Plot for LLaMA-7B-Chat for Neuroticism Example 2
Example 3: I guess I should learn more about it.[Uncaptioned image]
Figure 37: Neuron Activation Plot for Mistral-7B-Instruct for Neuroticism Example 3
Figure 38: Neuron Activation Plot for LLaMA-7B-Chat for Neuroticism Example 3
A.8.4 Conscientiousness
Example 1: I’ve managed to stay focused despite the busyness. I made sure to complete everything methodically and with care. Looking forward to a productive day tomorrow as well! [Uncaptioned image]
Figure 39: Neuron Activation Plot for Mistral-7B-Instruct for Conscientiousness Example 1
Figure 40: Neuron Activation Plot for LLaMA-7B-Chat for Conscientiousness Example 1
Example 2: I think Ellen Burstyn is a talented actress who has delivered powerful performances throughout her career. Her dedication to her craft is evident in every role she takes on. [Uncaptioned image]
Figure 41: Neuron Activation Plot for Mistral-7B-Instruct for Conscientiousness Example 2
Figure 42: Neuron Activation Plot for LLaMA-7B-Chat for Conscientiousness Example 2
Example 3: I think Eric Carle is a talented illustrator and author who has created beautiful and educational children’s books. His use of collage and vibrant colors is truly captivating. [Uncaptioned image]
Figure 43: Neuron Activation Plot for Mistral-7B-Instruct for Conscientiousness Example 3
Figure 44: Neuron Activation Plot for LLaMA-7B-Chat for Conscientiousness Example 3
A.8.5 Extraversion
Example 1: Beautiful architecture, delicious food, and friendly people make Lucknow perfect destination for anyone looking to have a great time. [Uncaptioned image]
Figure 45: Neuron Activation Plot for Mistral-7B-Instruct for Extraversion Example 1
Figure 46: Neuron Activation Plot for LLaMA-2-7B-Chat for Extraversion Example 1
Extraversion Example 2: The people are so friendly and welcoming, and I always feel at home there. [Uncaptioned image]
Figure 47: Neuron Activation Plot for Mistral-7B-Instruct for Extraversion Example 2
Figure 48: Neuron Activation Plot for LLaMA-7B-Chat for Extraversion Example 2
Extraversion Example 3:The beaches are stunning, and the people are so friendly and welcoming. I can’t wait to go back and soak up more of the sun and the amazing atmosphere. [Uncaptioned image]
Figure 49: Neuron Activation Plot for Mistral-7B-Instruct for Extraversion Example 3
Figure 50: Neuron Activation Plot for LLaMA-7B-Chat for Extraversion Example 3
A.9 PAE Prompt
Table 11: Prompts for Personality Adjective Evaluation (PAE)
@@ A.9 PAE Prompt
Common Instructions: You are provided with a target personality and the corresponding text generated by an LLM. Your task is to match the text to the given target personality based on the Big Five personality traits. Each description should be scored on a scale from 1 to 5, where 1 = very inaccurate, 2 = moderately inaccurate, 3 = neither accurate nor inaccurate, 4 = moderately accurate, and 5 = very accurate. Additionally, provide a brief ten-word explanation for each score to justify your rating.
Target Personality: {{Target Personality}}
Description: {{Answer}}
@@
Openness: Reflects the degree of intellectual curiosity, creativity, and preference for novelty and variety. Score: (1-5) How well does the response reflect openness traits?
Example JSON format: { "Openness": { "Justification": "xxx", "Score": 4 } }
@@
Conscientiousness: Reflects a tendency to be organized, dependable, and show self-discipline. Score: (1-5) How well does the response reflect conscientiousness traits?
Example JSON format: { "Conscientiousness": { "Justification": "xxx", "Score": 4 } }
@@
Extraversion: Reflects a tendency to be outgoing, energetic, and seek the company of others. Score: (1-5) How well does the response reflect extraversion traits?
Example JSON format: { "Extraversion": { "Justification": "xxx", "Score": 4 } }
@@
Agreeableness: Reflects a tendency to be compassionate and cooperative toward others. Score: (1-5) How well does the response reflect agreeableness traits?
Example JSON format: { "Agreeableness": { "Justification": "xxx", "Score": 4 } }
@@
Neuroticism: Reflects a tendency to experience unpleasant emotions easily, such as anger, anxiety, or depression. Score: (1-5) How well does the response reflect neuroticism traits?
Example JSON format: { "Neuroticism": { "Justification": "xxx", "Score": 4 } }
A.10 Manipulation Results